# Towards numerical two-loop integrand reduction

### v1

@@ 2.1 Basic notation and conventions
where $k_{1},\dots,k_{L}$ are loop momenta, $[dk_{i}]$ denotes the integration measure associated to the loop momentum $k_{i}$, and $\{p\}$ represents the set of external kinematics. The amplitude consists generally of $G$ partial contributions, individually denoted by $g$. Depending on the considered approach, the latter may be individual Feynman diagrams or consistently defined sub-amplitudes. In the following, we will refer to them as loop topologies. In what follows, we focus on the genuine two-loop contributions to the amplitude (so-called Theta topologies), namely those that cannot be trivially reduced to the product of one-loop ones (Infinity and Dumbbell topologies) [^1] For the terminology, see Figure 1 of Ref. Bevilacqua:2024fec .. Each loop topology is characterized by a numerator ${\cal N}$, which is a function of external momenta and wavefunctions as well as loop momenta, and by a product of $n$ inverse propagators $D_{1},\dots,D_{n}$, appearing in its denominator [^2] Inverse propagators raised in higher integer powers must also be included to accommodate special contributions..
@@
${\cal N}=P^{(n)}+\sum_{i=1}^{n}P_{i}^{(n-1)}D_{i}+\sum_{i=1}^{n-1}\sum_{j>i}^{n}P_{ij}^{(n-2)}D_{i}D_{j}+\ldots$ (4)
Plugging this expression in Eq. (1) results in a splitting of the integrand into terms with a different number of denominators. The quantities $P^{(n)}$, $P_{i}^{(n-1}$, $P_{ij}^{(n-2)},\dots$ appearing in Eq. (4) are polynomials of the form
@@ 2.2 Linear fit and fit by cut approach
where the coefficients $x_{i}$ and $y_{i}$ are expressible in terms of scalar products of the form $k_{i}\cdot p_{j}$. The latter coefficients characterize the loop momenta in $d=4$ dimensions. Complemented by $\mu_{11}$, $\mu_{12}$ and $\mu_{22}$ [^3] In $d=4-2\epsilon$, $\mu_{ij}\equiv k^{(\epsilon)}_{i}\cdot k^{(\epsilon)}_{j}$, with $k^{(\epsilon)}$ the components of the loop momenta beyond $d=4$ dimensions., they form a set of eleven variables which characterizes completely the loop momenta in $d=4-2\epsilon$ dimensions : $\{x_{1},x_{2},x_{3},x_{4},y_{1},y_{2},y_{3},y_{4},\mu_{11},\mu_{12},\mu_{22}\}$. We solve cut equations [^4] For alternative approaches, see Refs. Kosower:2011ty ; Mastrolia:2011pr ; Zhang:2012ce ; Badger:2012dp ; Badger:2013gxa ; Ita:2015tya . in terms of these variables, as we will see in section 3. In $d=4-2\epsilon$, the solution to the cut equations is unique in terms of the ISP, whereas in $d=4$ we usually have disjoint branches Zhang:2012ce , see section 3.1.1 for an explicit example.
@@
Returning to the solution of Eq. (4), let us first address the case of $d=4-2\epsilon$ dimensions. The cut equations fix a subset of the eleven parameters needed to fully describe the loop momenta. Assuming that the set of monomials $m_{i}$ $(i=1,\ldots,M)$ parametrizing a given polynomial $P$ is established, then an $M\times M$ matrix, $\cal M$, is obtained by evaluating the monomials on the solution to the cut equation, by assigning random values to its free parameters. The numerator can be cast in the form
${\cal N}={\cal N}_{4}+\sum\epsilon^{i}{\cal N}^{(i)}_{\epsilon}\,,$ (13)
where both ${\cal N}_{4}={\cal N}|_{d=4}$ and ${\cal N}^{(i)}_{\epsilon}$ are accessible numerically and explicitly depending on $\mu_{ij}$. These terms are used to calculate, with the same random values, the $M\times 1$ matrices ${\cal B}^{(4)}$, ${\cal B}^{(i)}$. Then the given polynomial $P$ is written explicitly as
$P=\sum_{i=1}^{M}\left(c_{i}^{(d=4)}+\sum\epsilon^{j}\;c_{i}^{({j})}\right)m_{i}$ (14)
@@
${\vec{c}\;}^{(d=4)}={\cal M}^{-1}{\cal B}^{(4)}\;\;\;\;\;{\vec{c}\;}^{({i})}={\cal M}^{-1}{\cal B}^{(i)}\;\;\;\;\;$ (15)
After the whole iterative procedure is completed, the so-called $N=N$ test is performed. The latter consists of checking the validity of Eq. (4) for arbitrary assignment of numerical values for all free parameters of the loop kinematics. We have checked that the $N=N$ test is fulfilled when the polynomials $P$ are constructed directly from the analytic expression of the numerator, as well as when using BasisDet to construct the ansatz for the polynomials.
In $d=4$ we obtain several disjoint solutions of the cut equations. On each branch, we have checked analytically that the $d=4$ numerator assumes a different form. On the other hand, the set of monomials obtained previously in $d=4-2\epsilon$ dimensions, contains linear dependencies due to the fact that in $d=4$ the two loop momenta and the three independent external momenta in a 4-particle amplitude for instance, is an over complete set and Gram determinants among them vanish, leading to non-trivial relations. In that case, the set of monomials $m_{i}$, provided by BasisDet, is evaluated at each branch of the cut-equation solution. Assuming the existence of $r$ branches, the matrix $\cal M$ has size $(rM)\times M$ and ${\cal B}_{0}^{(4)}$, sized $(rM)\times 1$, is calculated using ${\cal N}_{4,0}={\cal N}|_{\mu_{ij}=0,d=4}$. The system
@@
can still be solved with standard Linear Algebra algorithms such as QR decomposition, as long as the rank of the matrix is full, namely $rank({\cal M})=M$. As we will see later, this is true in most cases, but solutions can still be obtained in cases where the matrix is rank-deficient, $rank({\cal M})<M$. Notice that although in $d=4-2\epsilon$ case the reduction of the amplitude is complete, in $d=4$, the so-called rational terms need to be calculated Ossola:2008xq ; Badger:2008cm ; Pozzorini:2020hkx ; Lang:2020nnl ; Lang:2021hnw .
@@ 2.3 Global fit approach
Assuming that the numerator ${\cal N}$ is a function of the eleven (eight) parameters in $d=4-2\epsilon$ ($d=4$) dimensions, and a numerator provider (such as HELAC-2LOOP) is capable of returning numerical values for ${\cal N}_{4}$ and ${\cal N}_{\epsilon}$, we can instead determine all the coefficients of a single polynomial $P$ by means of a global fit Peraro:2019okx ; Badger:2021imn . In this method, all monomials involved in the polynomial $P$ are evaluated for several arbitrary random numerical assignments of the above-mentioned parameters. In this way, we construct a single square matrix ${\cal M}$, with size $M\times M$, where $M$ is the number of the monomials appearing in $P$. Using the same random values of the parameters, we also sample the numerator functions ${\cal N}_{4}$, ${\cal N}_{\epsilon}$, ${\cal N}_{4,0}$ and build $M\times 1$ matrices that we denote ${\cal B}^{(4)}$, ${\cal B}^{(\epsilon)}$ and ${\cal B}_{0}^{(4)}$ respectively. The difference to the fit by cut approach is that the kinematics are not constrained to obey cut equations. While the number $M$ of monomials is much larger, since it accommodates all monomials from all polynomials in Eq. (4), there are no significant conceptual differences between the two approaches. As we will show later, the solution can be obtained by using standard Linear Algebra packages, such as Eigen eigenweb and LAPACK LaPack , at both double and quadruple precision.
@@ 2.4 Projecting over a full family
will generically lead to polynomials $P$ depending on more variables rather than cutting over all 9 propagators,
@@
In the previous equations, $v_{i}\equiv k_{i}\cdot\eta$, where $\eta$ is the vector transverse to $p_{1},p_{2},p_{3}$ necessary to complete the 4-dimensional physical space, while $u_{I}$, $u^{i}_{I}$, $u^{i,j}_{I}$, $u^{i,j,k}_{I}$, $u^{i,j,k,l}_{I},\,\dots$, stand for ISP in the corresponding cut.
The same is true for a five-particle amplitude, where cutting over 8 propagators results in
@@ 3 Application of the method
In this section, the application of the method described in section 2 is demonstrated through several examples. The considered numerators are representative of a variety of scattering processes and consist of different kinematic dependencies, ranging from four-point to six-point kinematics. For the generation of the analytic expressions required, we used the Mathematica packages FeynArts Hahn:2000kx and FeynCalc Shtabovenko:2023idz , except for the 6-particle case, section 3.3.1, which has been generated by FORM Ruijl:2017dtg .
@@ 3.1.1 Double-box topologies
Above and henceforth, the shorthand notation $p_{i\ldots j}=p_{i}+\ldots+p_{j}$ will be used to denote the sum of the incoming external momenta, and $s=(p_{1}+p_{2})^{2}$ and $t=(p_{2}+p_{3})^{2}$ the standard Mandelstam variables. For all analyses presented below, we have used the following helicity assignment: $\lambda_{1}=+1,\lambda_{2}=+1,\lambda_{3}=-1,\lambda_{4}=-1$, for the helicities of the external gluons, all assumed to be incoming.
@@
    $\begin{gathered}k_{2}\cdot p_{2}\to\frac{s}{2}-k_{2}\cdot p_{1},\quad k_{1}\cdot p_{2}\to-\frac{s}{2},\quad k_{2}\cdot p_{3}\to-\frac{s}{2},\\ k_{1}\cdot k_{2}\to k_{1}\cdot p_{1},\quad k_{1}\cdot k_{1}\to-2k_{1}\cdot p_{1},\quad k_{2}\cdot k_{2}\to 0\,,\end{gathered}$ (28)
@@
    $\begin{gathered}k_{2}\cdot p_{1}\to 0,\quad k_{2}\cdot p_{2}\to\frac{s}{2},\quad k_{1}\cdot p_{2}\to-\frac{s}{2},\quad k_{2}\cdot p_{3}\to-\frac{s}{2},\quad k_{1}\cdot p_{3}\to\frac{s}{2},\\ k_{1}\cdot k_{2}\to k_{1}\cdot p_{1},\quad k_{1}\cdot k_{1}\to-2k_{1}\cdot p_{1},\quad k_{2}\cdot k_{2}\to 0\,,\end{gathered}$ (32)
@@
    For comparison, the cut conditions for 9 propagators are
@@
    with $r=\sqrt{-t(s+t)}$. In a numerical setup, for instance, calculating on the kinematic point $s=1,\,t=-1/5$, the 7-cut is represented by
@@
    We can now determine the coefficients of the polynomial $P_{7}$, which in this case has 70 coefficients, by calculating the numerator and the monomials of the basis, using 70 random assignments of the undetermined variables, $x_{3},x_{4},y_{3},y_{4}$, and solving the corresponding matrix equation
@@
    Table 3: Max Power for $k_{1}$, $k_{2}$ and $k_{1}$, $k_{2}$ combined, used as inputs for BasisDet.
@@ 3.1.2 Penta-triangle topologies
    This presents quite an interesting theoretical challenge for the construction of the ansatz: On the one hand, we have verified that choosing a loose ansatz, meaning a polynomial that leads to a rank deficient matrix is in fact not a problem, as long as the system is still solvable and the $N=N$ test is not affected. This may continue to hold if one chooses a polynomial basis which leads to an unsolvable linear equation, which we also found to be a potential pitfall during our study of this particular topology in $d=4$. On the other hand, restricting oneself to a smaller basis of polynomials, meaning polynomials containing a very small numer of monomials can lead to the following problem: While the solution to the linear equation may even be unique (i.e. full rank of the matrix ${\cal M}$), the functional behavior of the numerator may not be properly accounted for, leading to an incorrect solution which does not reproduce the $N=N$ test. The issue, therefore, of choosing an ansatz, i.e., a proper basis of polynomials to describe a given numerator is not at all trivial and still somewhat remains an open question, especially important if one aims to perform a $d=4$ fit. While BasisDet seems to be the best option available, it still runs the risk of falling into one of the two above problems, depending on the inputs given.
@@ 3.2.1 Penta-box topologies
The external kinematics is described by 4 independent momenta and five independent invariants, which can be chosen to be $S_{5}=\{s_{12},s_{23},s_{34},s_{45},s_{15}\}$, where $s_{ij}=(p_{i}+p_{j})^{2}$. In order to proceed with the reduction in an analytic setup, the expression of the numerator needs to be expressed in terms of the 11 invariants $V_{11}$, $k_{i}\cdot k_{j},$ with $i,j=1,2$, and $k_{i}\cdot p_{j}$ with $i=1,2,j=1,\ldots,4$. Since the analytic expression of the numerator involves the polarization vectors of the external gluons, $\varepsilon_{i},i=1,\ldots,5$, scalar products of the form $k_{i}\cdot\varepsilon_{j}$ and $p_{i}\cdot\varepsilon_{j}$ need to be expressed in terms of the variables $S_{5}$ and $V_{11}$. To this end, the following relation is used
@@ 3.2.2 $t\bar{t}H$
Table 13: $gg\to t\bar{t}H$ linear fit information beginning with 8-cut.
@@ 3.3.1 Six-gluon topology
and the reduction data are summarized in Tab. 15
@@ 4 Summary and Discussion
  • We have successfully studied the reduction equations for several cases in 4--, 5-- and 6--particle topologies, with massless and massive external particles and internal propagators: double-box, penta-triangle, hexa-bubble, non-planar double-box, $t\bar{t}$, penta-box, $t\bar{t}H$ and 6-gluon with a closed quark loop. We conclude that two-loop reduction at the integrand level is within reach for arbitrary processes. We plan to implement the solutions of the cut equations in a numerical setup Ossola:2007ax , in order to allow for a fully-fledged numerical reduction of two-loop integrand numerators.

### v3

@@ 2.1 Basic notation and conventions
where $k_{1},\dots,k_{L}$ are loop momenta, $dk_{i}]$ denotes the integration measure associated to the loop momentum $k_{i}$, and $\{p\}$ represents the set of external kinematics. The amplitude consists generally of $G$ partial contributions, individually denoted by $g$. Depending on the considered approach, the latter may be individual Feynman diagrams or consistently defined sub-amplitudes. In the following, we will refer to them as loop topologies. In what follows, we focus on the genuine two-loop contributions to the amplitude (so-called Theta topologies), namely those that cannot be trivially reduced to the product of one-loop ones (Infinity and Dumbbell topologies) [^1] For the terminology, see Figure 1 of Ref. Bevilacqua:2024fec .. Each loop topology is characterized by a numerator $\cal N$}, which is a function of external momenta and wavefunctions as well as loop momenta, and by a product of $n$ inverse propagators $D_{1},\dots,D_{n}$, appearing in its denominator [^2] Inverse propagators raised in higher integer powers must also be included to accommodate special contributions..
@@
${\cal N}=P^{(n)}+\sum_{i=1}^{n}P_{i}^{(n-1)}D_{i}+\sum_{i=1}^{n-1}\sum_{j>i}^{n}P_{ij}^{(n-2)}D_{i}D_{j}+\ldots+P_{12\ldots n}^{(0)}D_{1}D_{2}\ldots D_{n}$ (4)
The equation holds, as in the case of one loop Ossola:2006us , whenever the numerator of the amplitude has a rank, in $k_{1}$ and $k_{2}$, that does not exceed the number of corresponding propagators. In fact, as we will see later, the expansion may end long before its final contribution, depending on the rank of the numerator in question (see also Eq. (22) and section 3). Plugging this expression in Eq. (1) results in a splitting of the integrand into terms with a different number of denominators. The quantities $P^{(n)}$, $P_{i}^{(n-1)}$, $P_{ij}^{(n-2)},\dots$ appearing in Eq. (4) are polynomials of the form
@@ 2.2 Linear fit and fit by cut approach
where the coefficients $x_{i}$ and $y_{i}$ are expressible in terms of scalar products of the form $k_{i}\cdot p_{j}$. The latter coefficients characterize the loop momenta in $d=4$ dimensions. Complemented by $\mu_{11}$, $\mu_{12}$ and $\mu_{22}$ [^3] In $d=4-2\epsilon$, $\mu_{ij}\equiv k^{(\epsilon)}_{i}\cdot k^{(\epsilon)}_{j}$, with $k^{(\epsilon)}$ the components of the loop momenta beyond $d=4$ dimensions., they form a set of eleven variables which characterizes completely the loop momenta in $d=4-2\epsilon$ dimensions: $\vec{X}=\{x_{1},x_{2},x_{3},x_{4},y_{1},y_{2},y_{3},y_{4},\mu_{11},\mu_{12},\mu_{22}\}$. We solve cut equations [^4] For alternative approaches, see Refs. Kosower:2011ty ; Mastrolia:2011pr ; Zhang:2012ce ; Badger:2012dp ; Badger:2013gxa ; Ita:2015tya . in terms of these variables, as we will see in section˜3. In $d=4-2\epsilon$, the solution to the cut equations is unique in terms of the ISP, whereas in $d=4$ we usually have disjoint branches Zhang:2012ce , see section˜3.1.1 for an explicit example.
@@
[…]
${\cal N}\equiv{\cal N}(\vec{X},d)={\cal N}_{0}+\sum_{i\geq 1}\epsilon^{i}{\cal N}^{(i)}_{\epsilon}\,$, (13)
by expanding in powers of $\epsilon\equiv(4-d)/2$, where both ${\cal N}_{0}\equiv{\cal N}(\vec{X},d)|_{d=4}$ and ${\cal N}^{(i)}_{\epsilon}\equiv{\cal N}^{(i)}_{\epsilon}(\vec{X})$, are accessible numerically and depending on $\mu_{ij}$ through $\vec{X}$. These terms are used to calculate the $M\times 1$ matrices, ${\cal B}^{(0)}_{j}={\cal N}_{0}(\vec{X}_{j})$, ${\cal B}^{(i)}_{j}={\cal N}^{(i)}_{\epsilon}(\vec{X}_{j})$. Then the given polynomial $P$ is written explicitly as
$P=\sum_{i=1}^{M}\left(c_{i}^{(0)}+\sum\epsilon^{j}\;c_{i}^{({j})}\right)m_{i}$ (14)
@@
${\vec{c}\;}^{(0)}={\cal M}^{-1}{\cal B}^{(0)}\;\;\;\;\;{\vec{c}\;}^{({i})}={\cal M}^{-1}{\cal B}^{(i)}\;\;\;\;\;$ (15)
After the whole iterative procedure is completed, the so-called $N=N$ test is performed. The latter consists of checking the validity of Eq. (4) for arbitrary assignment of numerical values for all free parameters of the loop kinematics, $\vec{X}$, not restricted by cut equations. We have checked that the $N=N$ test is fulfilled when the polynomials $P$ are constructed directly from the analytic expression of the numerator, as well as when using BasisDet to construct the ansatz for the polynomials.
[…]
@@
can still be solved with standard Linear Algebra algorithms such as QR decomposition, as long as the rank of the matrix is full, namely $rank({\cal M})=M$. As we will see later, this is true in most cases, but solutions can still be obtained in cases where the matrix is rank-deficient, $rank({\cal M})<M$. Notice that in $d=4-2\epsilon$ case, when the information on the dependence of the numerator on $d$ and $\mu_{ij}$ is available, the reduction of the amplitude is complete, whereas in four dimensions, where this information is not available, the so-called rational terms need to be calculated in addition Ossola:2008xq ; Badger:2008cm ; Pozzorini:2020hkx ; Lang:2020nnl ; Lang:2021hnw .
@@ 2.3 Global fit approach
[…]
@@ 2.4 Projecting over a full family
with $v_{i}\equiv k_{i}\cdot\eta$, where $\eta$ is the vector transverse to $p_{1},p_{2},p_{3}$ necessary to complete the 4-dimensional physical space, while $u^{i}_{I}$, $u^{i,j}_{I}$, $u^{i,j,k}_{I}$, $u^{i,j,k,l}_{I},\,\dots$, stand for ISP in the corresponding cut, will generically lead to polynomials $P$ depending on more variables rather than cutting over all 9 propagators,
@@
where now $u_{i}$ stand for ISP in the corresponding cut (see explicit expressions in section 3).
The same is true for a five-particle amplitude, where now the physical phase-space is fully described by the four external independent momenta, cutting over 8 propagators results in
@@ 3 Application of the method
In this section, the application of the method described in section˜2 is demonstrated through several examples. The considered numerators are representative of a variety of scattering processes and consist of different kinematic dependencies, ranging from four-point to six-point kinematics. For the generation of the analytic expressions required, we used the Mathematica packages FeynArts Hahn:2000kx and FeynCalc Shtabovenko:2023idz , except for the 6-particle case, section 3.3.1, which has been generated by FORM Ruijl:2017dtg .
@@ 3.1.1 Double-box topologies
Above and henceforth, the shorthand notation $p_{i\ldots j}=p_{i}+\ldots+p_{j}$ will be used to denote the sum of the incoming external on-shell momenta, and $s=(p_{1}+p_{2})^{2}$ and $t=(p_{2}+p_{3})^{2}$ the standard Mandelstam variables. For all analyses presented below we have used the following helicity assignment: $\lambda_{1}=+1,\lambda_{2}=+1,\lambda_{3}=-1,\lambda_{4}=-1$, for the helicities of the external gluons, all assumed to be incoming.
@@
[…]
@@
[…]
@@
    For comparison, the cut conditions for 9 propagators, Eq. (30), are
@@
    with $r=\sqrt{-t(s+t)}$. In a numerical setup, for instance, calculating on the kinematic point s=1,t=1/5s=1,\,t=-1/5, the 7-cut is represented by
@@
    We can now determine the coefficients of the polynomial $P_{7}$, which in this case has 70 coefficients, by calculating the numerator $\cal N$}, and the monomials $m_{i}$ of the basis, using 70 random assignments of the undetermined variables, $x_{3},x_{4},y_{3},y_{4}$, i.e. ${\cal N}_{i}$ and $M_{ij}$ respectively, and solving the corresponding matrix equation
@@
    Table 3: Input used for BasisDet, denoting by $\{\{1,0\},n_{1}\}$ the maximal power $n_{1}$ in $k_{1}$, by $\{\{0,1\},n_{2}\}$ the maximal power $n_{2}$ in $k_{2}$ and by $\{\{1,1\},n_{12}\}$ the maximal power $n_{12}$ in $k_{1}$, $k_{2}$ combined.
@@ 3.1.2 Penta-triangle topologies
[…]
@@ 3.2.1 Penta-box topologies
The external kinematics is described by 4 independent momenta and five independent invariants, which can be chosen to be $S_{5}=\{s_{12},s_{23},s_{34},s_{45},s_{15}\}$, where $s_{ij}=(p_{i}+p_{j})^{2}$. In order to proceed with the reduction in an analytic setup, the expression of the numerator needs to be expressed in terms of the 11 invariants $V_{11}$, defined as the set of variables $k_{i}\cdot k_{j},$ with $i,j=1,2$, and $k_{i}\cdot p_{j}$ with $i=1,2,\;j=1,\ldots,4$. Since the analytic expression of the numerator involves the polarization vectors of the external gluons, $\varepsilon_{i},i=1,\ldots,5$, scalar products of the form $k_{i}\cdot\varepsilon_{j}$ and $p_{i}\cdot\varepsilon_{j}$ need to be expressed in terms of the set of variables $S_{5}$ and $V_{11}$. To this end, the following relation is used
@@ 3.2.2 $t\bar{t}H$
Table 13: $gg\to t\bar{t}H$ linear fit information beginning with 11-cut.
@@ 3.3.1 Six-gluon topology
and the reduction data are summarized in Tab. 15.
@@ 4 Summary and Discussion
  • We have successfully studied the reduction equations for several cases in 4-, 5- and 6-particle topologies, with massless and massive external particles and internal propagators: double-box, penta-triangle, hexa-bubble, non-planar double-box, $t\bar{t}$, penta-box, $t\bar{t}H$ and 6-gluon with a closed quark loop. We conclude that two-loop reduction at the integrand level is within reach for arbitrary processes. We plan to implement the solutions of the cut equations in a numerical setup Ossola:2007ax , in order to allow for a fully-fledged numerical reduction of two-loop integrand numerators.